OpenStudy (anonymous):

3 sqrt of 125x^3-2 sqrt of 45x^3

OpenStudy (anonymous):

possible answers are 6^15 6 · 3^15 6 · 3^14

OpenStudy (anonymous):

\[3\sqrt{125x ^{3}}-2\sqrt{45x ^{3}}=15x \sqrt{3x}-6x \sqrt{3x}=9x \sqrt{3x}\]

OpenStudy (anonymous):

mr.joselson is my answer 6 times 3 to the 15th power>?

OpenStudy (anonymous):

I am not sure of the question then

OpenStudy (anonymous):

sure of the what? my computer is messed up and the good answer thing is covering up what you said you werent sure of so can you jus retype what part your not sure of?

OpenStudy (anonymous):

Is there a variable x in the problem?

OpenStudy (anonymous):

a variable in where?

OpenStudy (anonymous):

Is the equation I wrote the same as the question?

OpenStudy (anonymous):

does sqrt refer to the numbers only or to whole terms in x^3

OpenStudy (anonymous):

only with the ^

OpenStudy (anonymous):

9 sqrt(5) x^3

OpenStudy (anonymous):

thats not even an option.options are at the top go take a look.

OpenStudy (anonymous):

I know, that's what your equation says though.

OpenStudy (anonymous):

expression not equation

OpenStudy (anonymous):

what yo0u mean? like you solved it and thats what you got?

OpenStudy (anonymous):

I simplified your expression

OpenStudy (anonymous):

ohh mr.i forgot your name yes its wrong how you wrote its too long

OpenStudy (anonymous):

estudier go to the top and see how he wrote it it stops before the = sign!!!! as in the first are right the rest it idk

OpenStudy (anonymous):

Your options have no x in them. The answer to your expression must include an x.

OpenStudy (anonymous):

Here is what I see as the "question": 3 sqrt of 125x^3-2 sqrt of 45x^3

OpenStudy (anonymous):

yeahh it does have an x in it sorry i was lookin ata completly diff problem

OpenStudy (anonymous):

your right sorry!!!

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

mr.jos can you go view what i jus posted